# Hyperons during proto-neutron star deleptonization and the emission of dark flavored particles

### v1

@@ Abstract
Complementary to high-energy experimental efforts, indirect astrophysical searches of particles beyond the standard model have long been pursued. The present article follows the latter approach and considers, for the first time, the self-consistent treatment of the energy losses from dark flavored particles produced in the decay of hyperons during a core-collapse supernova (CCSN). To this end, general relativistic supernova simulations in spherical symmetry are performed, featuring six-species Boltzmann neutrino transport, and covering the long-term evolution of the nascent remnant proto-neutron star (PNS) deleptonization for several tens of seconds. A well-calibrated hyperon equation of state (EOS) is therefore implemented into the supernova simulations and tested against the corresponding nucleonic model. It is found that supernova observables, such as the neutrino signal, are robustly insensitive to the appearance of hyperons for the simulation times considered in the present study. The presence of hyperons enables an additional channel for the appearance of dark sector particles, which is considered at the level of the $\Lambda$ hyperon decay. Assuming massless particles that escape the PNS after being produced, these channels expedite the deleptonizing PNS and the cooling behaviour. This, in turn, shortens the neutrino emission timescale. The present study confirms the previously estimated upper limits on the corresponding branching ratios for low and high mass PNS, by effectively reducing the neutrino emission timescale by a factor of two. This is consistent with the classical argument deduced from the neutrino detection associated with SN1987A.
@@ 1 Introduction
Massive stars end their lifes in the event of a CCSN, when the stellar core collapses due to pressure losses from electron captures on protons bound in iron-group nuclei and the photodissociation of nuclei. The core collapse halts when the density exceeds the normal nuclear density, and the stellar core bounces back with the formation of a shock wave. The initial propagation of this bounce shock leads to the release of a $\nu_{e}$-burst when reaching the neutrinospheres of the last inelastic scattering. The associated energy loss, in combination with the energy loss due to the photodissociation of the still infalling matter onto the expanding dynamic bounce shock, causes the latter to stall and turn into an accretion shock. Revival of this standing accretion shock is considered the CCSN explosion engine (for recent reviews, c.f. Refs. [1, 2] and references therein).
@@
Most of the gravitational binding energy gain is stored in the central compact PNS, in the magnetic-rotational energy [3], as well as in thermal degrees of freedom and neutrinos of all flavours, on the order of several $10^{53}$ erg. Hence, the neutrino-heating mechanism for shock revival, first proposed in Ref. [4], has long been studied in modern CCSN simulations. These are based in multi-dimensional neutrino radiation hydrodynamics featuring neutrino transport, where in the presence of multi-dimensional hydrodynamics phenomena, such as convection and hydrodynamic instabilities, the neutrino-heating efficiency increases, compared to spherically symmetric CCSN models, which therefore generally fail to yield CCSN explosions. There are two exceptions, (i) low-mass stellar progenitors in the zero-age main sequence (ZAMS) mass range of 8–10 M [5, 6, 7] as well as ultra-stripped stellar progenitors that result from binary evolution [8], and (ii) CCSN explosions triggered due to a sufficiently strong first-order phase transition from normal nuclear matter to the quark-gluon plasma at high baryon density (c.f. Refs. [9, 10, 11, 12, 13, 14] and references therein).
The latter CCSN explosion mechanism relates to one of the largest uncertainties in CCSN modelling, namely, the state of matter under CCSN conditions, i.e. at high baryon density, in excess of saturation density, large isospin asymmetry given by the hadronic charge density (the proton abundance in the absence of other charged hadronic resonances, on the order of $Y_{p}=0.05-0.6$) and temperatures on the order of several tens of MeV (for a recent review of the role of the EOS in simulations of CCSN, see Ref. [15]). Besides the QCD deconfinement phase transition, which has lately been explored extensively also in the context of binary neutron star mergers [16, 17], the impact of hadronic degrees of freedom has been studied in the context of failed CCSN in Refs. [18, 19, 20], even though the potential impact of strange, as well as non-strange hadronic degrees of freedom has long been studied for neutron stars (see recent reviews [21, 22, 23, 24] and references therein.) Thereby, particular emphasis has been placed on the hyperon puzzle, which is the related to the softening of the supersaturation density EOS due to the presence of additional, heavy hadronic degrees of freedom, potentially violating the maximum neutron star mass constraint of about $2$ M, the latter was derived from observations of massive pulsars [25, 26, 27, 28, 27, 29]. Several possible solutions have been widely discussed in the literature. Among them, it has been put forward an early onset of the quark-hadron deconfinement phase transition at densities below which strange hadronic resonances would appear, or the existence of much more repulsive hadronic interactions at high baryon densities. However, hyperon-nucleon and hyperon-hyperon interactions are less understood than nucleon-nucleon interactions, due to the sparse knowledge of scattering data. This makes their potentials poorly constrained and dynamics far from understood (c.f. Ref. [30] and references therein). The theoretical understanding has been brought forward solve the problem of thermal production of hyperons at low baryon density for heavy-ion physics based on the coupled channel approach (c.f. Refs. [31, 32, 33] and references therein) and furthermore, final-state interaction analyses [34, 35, 36, 37] as well as femtoscopy studies (see Ref. [38] for a review and references therein) have become available as well as data on hypernuclear structure [39, 40, 41].
@@
Due to the poorly known hyperon EOS at high baryon density, effective hyperonic model EOS have long been employed in astrophysical studies, e.g., based on the relativistic mean-field approach that treats the unknown interactions via baryon-meson couplings. The present article implements the FSU2H hyperonic relativistic mean field model in simulations of CCSN, focusing on the PNS deleptonization phase, i.e. the evolution after the CCSN explosion onset has been triggered during which the nascent PNS deleptonizes via the emission of neutrinos of all flavors on a timescale of 10 seconds. This phase is ideal for studying the impact of hadronic physics as the central density increases continuously as a consequence of the deleptonization.
@@
The requirements for multi-purpose EOS for applications of CCSN studies is to cover a large parameter space. Therefore, the relativistic mean field (RMF) framework has long been developed (for recent reviews, see Refs. [42, 43] and references therein). Recent reviews of the RMF framework for hyperon EOS can be fond in Ref. [44, 23] [^1] Note the existence of the online service CompOSE repository that provides data tables for different state of the art EOS ready for use in astrophysical applications, nuclear physics and beyond [45, 46]..
@@
In addition to the impact that hyperons can potentially have in CCSN through the EOS, they can open potentially new cooling channels with the emission of novel, yet hypothetical, particles from the interior of the PNS. In particular, if the net energy carried away by these particles is comparable to that of neutrinos, then significant changes in the neutrino signal are predicted. These can be tested using the neutrino flux that was observed from SN1987A  [47, 48]. Such phenomenology has been studied extensively in the context of the QCD axions, produced in the stellar plasma by, e.g. nucleon-nucleon bremsstrahlung [49, 50, 51, 52, 53, 54, 55], from pions studied [56] and in simulations [57], and has been extended to many other “dark-sector” theories predicting the existence of new weakly interacting particles with mass $\lesssim\mathcal{O}(T_{\rm PNS})$ [58, 59, 60, 61, 62, 63].
@@
Adding “standard” degrees of freedom in the plasma, on top of nucleons and electrons, can also extend these analyses, as recently demonstrated for muons (c.f. Refs. [64, 65, 66, 67, 63] and references therein), for pions [68, 69, 70, 56, 57, 71], and for hyperons [72, 73, 74, 75]. In the case of hyperons, this was studied for various models that can induce flavor-changing neutral currents, such as the following decays $\Lambda\to n$ and $\Lambda\to\gamma$, where the energy difference between final and initial states is considered to be carried away by dark bosonic degree of freedom. Therefore, results of spherically symmetric simulations were employed without exotic cooling and implementing conventional EOS without hyperons. To establish limits on these models, the impact of the new dark emission on the dynamics of the CCSN was neglected, and thermodynamic quantities such as chemical potentials were re-derived by using interpolation tables of the hyperonic extensions of the EOS used in the simulations (see also Ref. [75] for a slightly different approach).
@@
A dark sector coupled to quarks can have a rich flavor structure that is subject to strangeness-changing neutral currents (for reviews, see Refs. [76, 77]). One prototypical example of these dark flavored sectors is the QCD axion, with non-diagonal flavor couplings, known as “familon”. This was first predicted in Ref. [78] as a consequence of jointly solving the strong CP and flavor problems, and was further developed [79, 80, 81, 72, 82]. Other examples include the massless dark photon, which interacts with standard-model fermions only through non-renormalizable operators (see Refs. [83, 84, 85] and references therein) or dark baryons [74]. The latter are neutral fermions with masses of $\approx 1$ GeV endowed with baryon number and that can be part of dark matter, and explain baryogenesis (c.f. Refs. [86, 87]) and the neutron lifetime puzzle [88]. The self-consistent implementation of hyperons in CCSN simulations enables us to overcome previous limitations (c.f. Ref. [73] and references therein), studying the feedback of dark photon production on the supernova dynamics and neutrino emission.
@@
The paper is organized as follows. In Sec. 2 the RMF framework for reference nucleonic and hyperonic EOS will be introduced, which is evaluated at selected conditions in Sec. 3, including finite temperature, and discussed accordingly. CCSN simulations will be launched and evaluated in Sec. 4. The comprehensive analysis from the impact of dark photons will be provided in Sec. 5 and the manuscript closes with a summary in Sec. 6.
2 Hyprons within the relativistic mean field framework
@@
The present study of the impact of strange degrees of freedom in CCSN simulations employed the FSU2H RMF EOS. In this framework, the interactions between the baryons is mediated via the exchange of virtual mesons, based on the following Lagrangian for baryons $\mathcal{L}_{\rm B}$ and mesons $\mathcal{L}_{\rm m}$,
@@
Table 1: Parameters of the model FSU2H. The mass of the nucleon is equal to $m_{N}=939$ MeV.
$m_{\sigma}$ $m_{\omega}$ $m_{\rho}$ $m_{\sigma^{*}}$ $m_{\phi}$ $g_{\sigma N}^{2}$ $g_{\omega N}^{2}$ $g_{\rho N}^{2}$ $\kappa$ $\lambda$ $\zeta$ $\Lambda_{\omega}$
[[[[MeV]]]] [[[[MeV]]]] [[[[MeV]]]] [[[[MeV]]]] [[[[MeV]]]] [[[[MeV]]]]
497.479 782.5 763 980 1020 102.72 169.53 197.27 4.0014 -0.0133 0.008 0.045
@@
where the expectation values of the scalar mesons, denoted as $\bar{\sigma}$ and $\bar{\sigma}^{*}$, govern their behaviour, while those of the vector mesons, i.e. $\bar{\omega}$, $\bar{\rho}$ and $\bar{\phi}$, control the different chemical potentials as follows,
@@ 3 Equation of state with hyperons
The different parameters of the FSU2H model are chosen so as to predict an EOS that is compatible with the constraints coming from heavy ion collisions as well as those obtained from the saturation properties of nuclear matter and finite nuclei. At the same time the model predicts a maximum neutron star mass that is above 2.0 M and a radius about 13 km for canonical neutron stars of 1.4 M (for details, see Refs. [89, 90]). This model was then extended to finite temperature [91], whereas more recently the hyperonic uncertainties have been analyzed for the determination of the masses, radii, tidal deformabilities and moments of inertia, both at zero and finite temperatures [92]. Indeed, the FSU2H model is compatible with the tidal deformability extracted from the GW170817 event [93] as well as the NICER determinations on radii [94, 95, 96, 97]. The values of the parameters can be found in Tables 1 and 2.
@@
Figures 1 and 2 illustrate selected quantities of the NUCLEONS EOS (black lines) and HYPERONS EOS (blue lines) at certain conditions, at $T=0$ in Fig. 1 and at a constant entropy of $s=3~{}k_{\rm B}$ in Fig. 2, both assuming $\beta$-equilibrium. For both cases it becomes evident that HYPERONS EOS is substantially softer at high densities than NUCLEONS, illustrated by the pressure in the upper left panels of Figs. 1 and 2. This is a consequence from the lower neutron abundance since the EOS is dominated by the neutron degeneracy at the high densities explored here (see therefore the hadron abundances in the upper right panels and note the generally low abundances [^2] In the case of particles with baryon number equal one, abundances $Y_{i}$ and mass fractions $X_{i}$ are identical, defined through the partial densities $n_{i}=Y_{i}\,n_{\rm B}$ with baryon density $n_{\rm B}$. of hyperons $\Lambda$, $\Sigma^{-}$ and $\Xi^{-}$). Note also the modified charge neutrality conditions in the presence of hyperons, illustrated via the electron and muon abundances, $Y_{e}$ and $Y_{\mu}$, respectively, in the lower right panels, together with the density dependence of the effective masses gap equations (2.4).
Figure 1:  Equation of state at $\beta$-equilibrium at $T=0$, comparing the HYPERON (blue lines) and the reference NUCLEON EOS (black lines), showing (from top left to bottom right) the total pressure $P$, selected mass fractions $X_{i}$, corresponding chemical potentials $\mu_{i}$, without restmass $m_{i}$ and the effective masses $m_{i}^{*}$ as well as the abundances of electrons and muons, $Y_{e}$ and $Y_{\mu}$, respectively.
@@
The assumption of $\beta$-equilibrium results in the condition of equal chemical potentials for electrons and muons, $\mu_{e}=\mu_{\mu}$, once the electron chemical potential exceeds the muon restmass, $\mu_{e}\geq m_{\mu}=106$ MeV, which was employed here for the calculation of the muon abundances. Note that this assumption differs from the conditions realized in the CCSN simulations. This discrepancy has been shown in Ref. [12] that while $\mu_{\mu}\rightarrow\mu_{e}$, this condition can be reached only after more than 30 s post bounce in simulations of the PNS deleptonization and cooling, studied based on the DD2 relativistic mean field EOS, with density dependent meson-nucleon couplings [98]. As a consequence of this simplification applied here for the calculation of the EOS, the muon abundance is found to be significantly larger than what we find in the CCSN and PNS simulations that will be discussed in Sec. 4 below.
The lower right panels in Figs. 1 and 2 also show the effective hadron masses, $m_{i}^{*}$ relative to their vacuum values $m_{i}$, from which it becomes evident that the neutron mass decreases most rapidly towards about 15% of its vacuum value at a density of slightly above $\rho=10^{15}$ g cm-3, whereas the effective mass of the $\Lambda$ is about 50% of its vacuum value at the same density. The corresponding chemical potentials are shown in the lower left panels of Figs. 1 and 2. At zero temperature, hyperons are strictly suppressed when their effective chemical potentials, shown in the lower left panel in Fig. 1, are below their respective restmasses, while at finite temperature, shown in Fig. 2, this condition does not apply any longer.

### v2

# Hyperons during proto-neutron star deleptonization and the emission of dark flavoured particles

@@ Abstract
Complementary to high-energy experimental efforts, indirect astrophysical searches of particles beyond the standard model have long been pursued. The present article follows the latter approach and considers, for the first time, the self-consistent treatment of the energy losses from dark flavoured particles produced in the decay of hyperons during a core-collapse supernova (CCSN). To this end, general relativistic supernova simulations in spherical symmetry are performed, featuring six-species Boltzmann neutrino transport, and covering the long-term evolution of the nascent remnant proto-neutron star (PNS) deleptonization for several tens of seconds. A well-calibrated hyperon equation of state (EOS) is therefore implemented into the supernova simulations and tested against the corresponding nucleonic model. It is found that supernova observables, such as the neutrino signal, are robustly insensitive to the appearance of hyperons for the simulation times considered in the present study. The presence of hyperons enables an additional channel for the appearance of dark sector particles, which is considered at the level of the $\Lambda$ hyperon decay. Assuming massless particles that escape the PNS after being produced, these channels expedite the deleptonizing PNS and the cooling behaviour. This, in turn, shortens the neutrino emission timescale. The present study confirms the previously estimated upper limits on the corresponding branching ratios for low and high mass PNS, by effectively reducing the neutrino emission timescale by a factor of two. This is consistent with the classical argument deduced from the neutrino detection associated with SN1987A.
@@ 1 Introduction
Massive stars end their lifes in the event of a CCSN, when the stellar core collapses due to pressure losses from electron captures on protons bound in iron-group nuclei and the photodissociation of nuclei. The core collapse halts when the density exceeds the normal nuclear density, and the stellar core bounces back with the formation of a shock wave. The initial propagation of this bounce shock leads to the release of a $\nu_{e}$-burst when reaching the neutrinospheres of the last inelastic scattering. The associated energy loss, in combination with the energy loss due to the photodissociation of the still infalling matter onto the expanding dynamic bounce shock, causes the latter to stall and turn into an accretion shock. Revival of this standing accretion shock is considered the CCSN explosion engine (for recent reviews, c.f., Refs. [1, 2, 3] and references therein).
@@
Most of the gravitational binding energy gain is stored in the central compact PNS, in the magnetic-rotational energy [4], as well as in thermal degrees of freedom and neutrinos of all flavours, on the order of several $10^{53}$ erg. Hence, the neutrino-heating mechanism for shock revival, first proposed in ref. [5], has long been studied in modern CCSN simulations. These are based in multi-dimensional neutrino radiation hydrodynamics featuring neutrino transport, where in the presence of multi-dimensional hydrodynamics phenomena, such as convection and hydrodynamic instabilities, the neutrino-heating efficiency increases, compared to spherically symmetric CCSN models, which therefore generally fail to yield CCSN explosions. There are two exceptions, (i) low-mass stellar progenitors in the zero-age main sequence (ZAMS) mass range of 8–10 M [6, 7, 8] as well as ultra-stripped stellar progenitors that result from binary evolution [9, 10], despite substantial differences between spherically symmetric and multi-dimensional simulations [11, 12, 13, 14, 15], and (ii) CCSN explosions triggered due to a sufficiently strong first-order phase transition from normal nuclear matter to the quark-gluon plasma at high baryon density (c.f. Refs. [16, 17, 18, 19, 20, 21] and references therein) as well as the spontaneous scalarization of beyond the standard model degrees of freedom [22].
[…]
@@
Due to the poorly known hyperon EOS at high baryon density, effective hyperonic model EOS have long been employed in astrophysical studies, e.g., based on the relativistic mean-field (RMF) approach that treats the unknown interactions via baryon-meson couplings. The present article implements the FSU2H hyperonic relativistic mean field model in simulations of CCSN, focusing on the PNS deleptonization phase, i.e. the evolution after the CCSN explosion onset has been triggered during which the nascent PNS deleptonizes via the emission of neutrinos of all flavours on a timescale of 10 seconds. This phase is ideal for studying the impact of hadronic physics as the central density increases continuously as a direct consequence of the deleptonization.
@@
The requirements for multi-purpose EOS for applications of CCSN studies is to cover a large parameter space. Therefore, the RMF framework has long been developed (for recent reviews, see Refs. [52, 53] and references therein). Recent reviews of the RMF framework for hyperon EOS can be fond in ref. [54, 33] [^1] Note the existence of the online service CompOSE repository that provides data tables for different state of the art EOS ready for use in astrophysical applications, nuclear physics and beyond [55, 56]..
@@
In addition to the impact that hyperons can potentially have in CCSN through the EOS, they can open potentially new cooling channels with the emission of novel, yet hypothetical, particles from the interior of the PNS. In particular, if the net energy carried away by these particles is comparable to that of neutrinos, then significant changes in the neutrino signal are predicted. These can be tested using the neutrino flux that was observed from SN1987A  [57, 58]. Such phenomenology has been studied extensively in the context of the QCD axions, produced in the stellar plasma by, e.g. nucleon-nucleon bremsstrahlung [59, 60, 61, 62, 63, 64, 65], from pions studied [66] and in simulations [67], and has been extended to many other “dark-sector” theories predicting the existence of new weakly interacting particles with mass $\lesssim\mathcal{O}(T_{\rm PNS})$ [68, 69, 70, 71, 72, 73], which is on the order of few tens of MeV.
@@
Adding “standard” degrees of freedom in the plasma, on top of nucleons and electrons/positrons, can also extend these analyses, as was recently demonstrated in the case of muons (c.f. Refs. [74, 75, 76, 77, 73] and references therein), for pions [78, 79, 80, 66, 67, 81], and for hyperons [82, 83, 84, 85]. In the case of hyperons, this was studied for various models that can induce flavour-changing neutral currents, such as the following decays, $\Lambda\to n$ and $\Lambda\to\gamma$, where the energy difference between final and initial states is considered to be carried away by dark degrees of freedom. Therefore, results of spherically symmetric simulations were employed without exotic cooling and implementing conventional EOS without hyperons. To establish limits on these models, the impact of the new dark emission on the dynamics of the CCSN was neglected, and thermodynamic quantities such as chemical potentials were re-derived by using interpolation tables of the hyperonic extensions of the EOS used in the simulations (see also ref. [85] for a slightly different approach).
@@
A dark sector coupled to quarks can have a rich flavour structure that is subject to strangeness-changing neutral currents (for reviews, see Refs. [86, 87]). One prototypical example of these dark flavoured sectors is the QCD axion, with non-diagonal flavour couplings, known as “familon”. This was first predicted in refs. [88, 89] as a consequence of jointly solving the strong CP and flavour problems, and was further developed [90, 91, 92, 82, 93]. Other examples include the dark photon, which interacts with standard-model fermions only through non-renormalizable operators (see Refs. [94, 95, 96, 97] and references therein) or dark baryons [84]. The latter are neutral fermions with masses of $\approx$1 GeV endowed with baryon number and that can be part of dark matter, and explain baryogenesis (c.f. Refs. [98, 99]) and the neutron lifetime puzzle [100]. The self-consistent implementation of hyperons in CCSN simulations enables us to overcome previous limitations (c.f. ref. [83] and references therein), studying the feedback of dark boson production on the supernova dynamics and neutrino emission.
@@
The paper is organized as follows. In section 2 the RMF framework for reference nucleonic and hyperonic EOS will be introduced, which is evaluated at selected conditions in section 3, including finite temperature, and discussed accordingly. CCSN simulations will be launched and evaluated in section 4. The comprehensive analysis from the impact of dark bosons will be provided in section 5 and the manuscript closes with a summary in section 6.
2 Hyperons within the relativistic mean field framework
@@
The present study of the impact of strange degrees of freedom in CCSN simulations employed the FSU2H RMF EOS. In this framework, the interactions between the baryons are mediated via the exchange of virtual mesons, based on the following Lagrangian for baryons $\mathcal{L}_{\rm B}$ and mesons $\mathcal{L}_{\rm m}$,
@@
Table 1: Parameters of the model FSU2H, meson masses $m_{i}$ and meson-nucleon coupling constants $g_{i}^{2}$. The mass of the nucleon is equal to $m_{N}=939$ MeV.
$m_{\sigma}$ $m_{\omega}$ $m_{\rho}$ $m_{\sigma^{*}}$ $m_{\phi}$ $g_{\sigma N}^{2}$ $g_{\omega N}^{2}$ $g_{\rho N}^{2}$
[[[[MeV]]]] [[[[MeV]]]] [[[[MeV]]]] [[[[MeV]]]] [[[[MeV]]]]
497.479 782.5 763 980 1020 102.72 169.53 197.27
Table 2: Other parameters of the Lagrangian (2.3) of the model FSU2H.
$\kappa$ $\lambda$ $\zeta$ $\Lambda_{\omega}$
[[[[MeV]]]]
4.0014 -0.0133 0.008 0.045
@@
where the expectation values of the scalar mesons, denoted as $\bar{\sigma}$ and $\bar{\sigma}^{*}$, govern their behaviour. Since no isospin splitting is considered due to the absence of the $\delta$ meson channel, $m_{N}=939$ MeV is used in FSU2H for the nucleon rest masses while those of the hyperons are given in the Particle Data Booklet [105]. The expectation values of the vector mesons, i.e. $\bar{\omega}$, $\bar{\rho}$ and $\bar{\phi}$, control the different chemical potentials as follows,
@@ 3 Equation of state with hyperons
The different parameters of the FSU2H model are chosen so as to predict an EOS that is compatible with the constraints coming from heavy ion collisions as well as those obtained from the saturation properties of nuclear matter and finite nuclei. At the same time the model predicts a maximum neutron star mass that is above 2.0 M and a radius about 13 km for canonical neutron stars of 1.4 M (for details, see Refs. [101, 102]). This model was then extended to finite temperature [103], whereas more recently the hyperonic uncertainties have been analyzed for the determination of the masses, radii, tidal deformabilities and moments of inertia, both at zero and finite temperatures [104]. Indeed, the FSU2H model is compatible with the tidal deformability extracted from the GW170817 event [106] as well as the NICER determinations on radii [107, 108, 109, 110]. The values of the parameter meson masses and meson-nucleon coupling constants can be found in Table 1, whereas the other parameters of the Lagrangian (2.3) are given in Table 2. The ratios of coupling constants of hyperons to mesons are give in Table 3.
@@
Figures 1 and 2 illustrate selected quantities of the NUCLEONS EOS (black lines) and HYPERONS EOS (blue lines) at certain conditions, at $T=0$ in Fig. 1 and at a constant entropy of $s=3~{}k_{\rm B}$ in Fig. 2, both assuming $\beta$-equilibrium. For both cases it becomes evident that the HYPERONS EOS is substantially softer at high densities than the NUCLEONS EOS, illustrated by the pressure in the upper left panels of Figs. 1 and 2. This is a consequence from the lower neutron abundance since the EOS is dominated by the neutron degeneracy at the high densities explored here (see therefore the hadron abundances in the upper right panels and note the generally low abundances222 In the case of particles with baryon number equal one, abundances $Y_{i}$ and mass fractions $X_{i}$ are identical, defined through the partial densities $n_{i}=Y_{i}\,n_{\rm B}$ with baryon density $n_{\rm B}$. of hyperons $\Lambda$, $\Sigma^{-}$ and $\Xi^{-}$). Note also the modified charge neutrality conditions in the presence of hyperons, illustrated via the electron and muon abundances, $Y_{e}$ and $Y_{\mu}$, respectively, in the lower right panels, together with the density dependence of the effective masses gap equations (2.4).
Figure 1:  Equation of state at $\beta$-equilibrium at $T=0$, comparing the HYPERON EOS(blue lines) and the reference NUCLEON EOS (black lines), showing (from top left to bottom right) the total pressure, $P$, selected mass fractions, $X_{i}$, corresponding chemical potentials, $\mu_{i}$ without restmass $m_{i}$ and the effective masses, $m_{i}^{*}$, as well as the abundances of electrons and muons, $Y_{e}$ and $Y_{\mu}$, respectively.
@@
The assumption of $\beta$-equilibrium results in the condition of equal chemical potentials for electrons and muons, $\mu_{e}=\mu_{\mu}$, once the electron chemical potential exceeds the muon restmass, $\mu_{e}\geq m_{\mu}=106$ MeV, which was employed here for the calculation of the muon abundances. Note that this assumption differs from the conditions realized in the CCSN simulations. This discrepancy has been pointed out in Refs. [118, 19], that while $\mu_{\mu}\rightarrow\mu_{e}$, this condition can be reached only after more than 30 s post bounce in simulations of the PNS deleptonization and cooling, studied based on the DD2 relativistic mean field EOS, with density dependent meson-nucleon couplings [111]. As a consequence of this simplification applied here for the calculation of the muon EOS, the muon abundance $Y_{\mu}$ is found to be significantly larger than what we find in the CCSN and PNS simulations, that will be discussed in Sec. 4 below.
The lower right panels in Figs. 1 and 2 also show the effective hadron masses, $m_{i}^{*}$, relative to their vacuum values $m_{i}$, from which it becomes evident that the neutron mass decreases most rapidly towards about 15% of its vacuum value at a density of slightly above $\rho=10^{15}$ g cm-3, whereas the effective mass of the $\Lambda$ is about 50% of its vacuum value at the same density. The corresponding chemical potentials, $\mu_{i}$ without rest masses $m_{i}$, are shown in the lower left panels of Figs. 1 and 2. At zero temperature, hyperons are strictly suppressed when their effective chemical potentials, shown in the lower left panel in Fig. 1, are below their respective rest masses, while at finite temperature, shown in Fig. 2, this condition does not apply any longer.